what is the exact area of the base of a circular swimming pool with diameter of 16ft?
i tried a=3.14xrxr x is the multiplication symbol
so 3.14 x 8 x 8 and it was equal to 200.96 which is none of the answer choices

Answer by lwsshak3(11628)

what is the exact area of the base of a circular swimming pool with diameter of 16ft?
i tried a=3.14xrxr x is the multiplication symbol
so 3.14 x 8 x 8 and it was equal to 200.96 which is none of the answer choices
**
I think the problem here is that they want the exact area so you must leave the answer in terms of π, because π is an irrational number. So this is what I believe is wanted:
Area=πr^2=π8^2=64π